OpenStudy (anonymous):

what is the geometric mean of 4 and 7

ganeshie8 (ganeshie8):

geometric mean of \(a\) and \(b\) = \(\sqrt{ab}\)

ganeshie8 (ganeshie8):

plugin the numbers...

OpenStudy (anonymous):

is the mean 4

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

mean = \(\sqrt{4 \times 7}\) = ?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

You would need to square 28 i believe though, the answer doesn't have to be a perfect square.

ganeshie8 (ganeshie8):

yes, you need to take square root of 28

ganeshie8 (ganeshie8):

mean = \(\sqrt{4 \times 7}\) = \(\sqrt{28}\) = ?

OpenStudy (anonymous):

another question a triangle has side of lengths 24,143 and 145 is it a right triangle

OpenStudy (anonymous):

thanks